# Composable Finite-Size Security of High-Dimensional Quantum Key Distribution Protocols

### v1

@@ V.2 Formalization of the protocol
For illustration purposes, we apply our method to a high-dimensional temporal entanglement setup using a Franson interferometer, analyzed in Ref. [27, Protocol 1]. A source prepares polarized photon pairs of the form |$\ket{\Psi_{1}}=\ket{\mathrm{DD}}\otimes\frac{1}{\sqrt{d}}\sum_{k=0}^{d-1}\ket{kk}$ which are distributed to the communicating parties. With probability $p_{1}$ each party performs a Time-of-Arrival (ToA) measurement, represented by $\{A_{1}^{x}\}_{x=0}^{d-1}$ and $\{B_{1}^{y}\}_{y=0}^{d-1}$, respectively, and with probability $1-p_{1}$ each of them performs a temporal superposition (TSUP) measurement of not necessarily neighboring time-bins, represented by $\{A_{2}^{x}\}_{x=0}^{d-1}$ and $\{B_{2}^{y}\}_{y=0}^{d-1}$, respectively. We start by deriving the protocol map representing the action of this protocol, specifying the map $\Phi$ introduced in Section III, following the formalism in [48, 49].
For each round, Alice and Bob may store a measurement outcome $\alpha_{i}$ and $\beta_{i}$ and make a public announcement $a_{i}$ and $b_{i}$, where the measurement outcomes are drawn from the sets $\mathcal{X}_{A}$ and $\mathcal{X}_{B}$ and the announcements from the set $\mathcal{A}_{A}$ and $\mathcal{A}_{B}$ for Alice and Bob, respectively. In our case, these announcements are the choice of measurement, i.e., a number $b$ related to the label of the POVM measurement performed. For the analyzed protocol, $b=1$ corresponds to time-of-arrival measurements, labeled ‘ToA’, while $b=2$ refers to superposition measurements of neighboring time bins, labeled ‘TSUP’. They store their measurement outcomes in their private registers $\bar{A}$ and $\bar{B}$ and their announcements in their public registers $\tilde{A}$ and $\tilde{B}$. Since we perform direct reconciliation, the key map $g$ saves the outcome of the map $g:~{}\mathcal{X}_{A}\times\mathcal{A}_{A}\times\mathcal{A}_{B}\rightarrow\{0,...,d-1,\perp\}$ to the key register $Z$. In more detail, since the protocol only generates key when both Alice and Bob measured ToA, the key map for our protocol is trivial and simply copies Alice’s measurement outcome contained in her private register $\bar{A}$ to the key register $Z$. This sets the stage for introducing the formal protocol map. Let $\rho_{AB}$ be Alice’s and Bob’s shared quantum state after the generation and distribution of the high-dimensionally entangled state and let $\rho_{ABE}$ denote Eve’s purification. Furthermore, since Alice’s and Bob’s announcements are public, Eve has access to them. We use the short notation $[E]$ to label Eve’s register including all available public information. Eve’s goal is to perform some (to us unknown) POVM measurement, aiming to gain as much information as possible about Alice’s measurement outcome. Since Alice and Bob only generate key if both measure the time of arrival, we obtain the following map, representing their measurement and announcement steps
@@
$\otimes\left[\left(\sqrt{A_{1}^{i}}\otimes\sqrt{B_{1}^{i}}\otimes\sqrt{E^{i}}\right)\rho_{ABE}\left(\sqrt{A_{1}^{i}}\otimes\sqrt{B_{1}^{i}}\otimes\sqrt{E^{i}}\right)^{\dagger}\right]$
@@
$\sigma_{Z[E]}$ $=\mathrm{Tr}_{A\bar{A}B\bar{B}}\left[V\Phi_{\mathrm{M\&A}}V^{\dagger}\right]$ (27)
However, since the announcements for this protocol are trivial, they do not give any additional information. Hence, after inserting the expressions for Alice’s and Bob’s key measurement POVM and tracing out the announcement registers, we obtain
@@ VI Results
In this section, we demonstrate our method by deriving composable finite-size key rates for the high-dimensional temporal entanglement protocol and setup analyzed in Ref. [27]. While this is one illustrative example, our method applies to general high-dimensional protocols in a very similar fashion.
In accordance with Ref. [31], we chose $d-1$ observables $\hat{W}_{k}:=q_{0}\sum_{i=0}^{d-1}\ket{i,i}\!\!\bra{i,i}+\sum_{z=1}^{d-1}\sum_{i=0}^{d-1}q_{z,i}\left(\ket{i,i}\!\!\bra{i+z,i+z}+\ket{i+z,i+z}\!\!\bra{i,i}\right)$ and $\hat{W}_{0}:=\sum_{\stackrel{{\scriptstyle i,j=0}}{{i\neq j}}}^{d-1}p\ket{i,j}\!\!\bra{i,j}$ |, where $p=1$. We chose the $q_{z,i}$ such that the $k$-th operator acts only on the $k$-th off-diagonal, which allows us to include the optimal choice of the relative weights between the witnesses in the optimization problem. More details regarding this choice and the intuition behind can be found in [31, Appendix D].
First, we cover the fixed-length scenario, which best suits theoretical discussions of achievable ‘one-shot’ key rates, followed by a discussion of variable-length key rates under specific noise models. In all examples given, we use $r_{S}$ to refer to the splitting ratio of the beam splitters in Alice’s and Bob’s labs that passively choose the measurement setting (ToA or TSUP). For example, a splitting ratio of $r_{S}=1\%$ means that $1\%$ of the photons are directed to the TSUP measurement apparatus, while the remaining $99\%$ are measured in ToA. Thus, they obtain $r_{S}^{2}N$ rounds of aligned TSUP measurements and $(1-r_{S})^{2}$ rounds of aligned ToA measurements. Given that $8(d-1)$ different click events are combined into $W_{1}$, we obtain $m_{W_{1}}:=\frac{r_{S}^{2}N}{8(d-1)}$. Additionally, we choose $m_{W_{2}}=m_{W_{1}}$ such that $n:=N(1-r_{S})^{2}-Nr_{s}^{2}=N(1-2r_{s})$, since the click elements required for $W_{2}$ are obtained from ToA measurements.
For demonstration purposes, if not mentioned otherwise, we assume an isotropic noise model $\rho=v\outerproduct{\Psi_{1}}{\Psi_{1}}+\frac{(1-v)}{d^{2}}\mathbbm{1}_{d^{2}}$. However, we want to emphasize that our method applies independently of the chosen noise model. Also, note that the leakage in the information reconciliation phase $\mathrm{leak}_{\mathrm{IR}}$ that appears in the key rate formulas given in Section III, is the sum of the error-correction leakage $\mathrm{leak}_{\mathrm{EC}}$ and the length of the hash used for error-verification $\log_{2}\left(\frac{2}{\epsilon_{\mathrm{EV}}}\right)$. If not stated otherwise, we assume ideal error-correction.
@@
[…]
@@ VI.1 Fixed-length key rates
We start with fixed-length keys, also called ‘one-shot’ key rates, corresponding to key rates usually reported in the literature. Recall, that for the definition of the acceptance set (see Theorem 1), one may choose to extend the set of accepted statistics by choosing the parameters $t_{X}>0$ (for each observable $X$). Intuitively, this extends the set of accepted statistics from a point to a (hyper-) rectangle with sides $t_{X}$ for $X\in\Theta$. To simplify notation and to describe the choices for the acceptance test with a single parameter, we measure the chosen $t_{X}$ for the different observables in multiples of their respective $\mu_{X}$ by introducing a factor $t_{F}$, $t_{X}=t_{F}\mu_{X}$. In the whole subsection discussing the results for the fixed-length scenario, we fix $t_{F}=1$. The choice of $t_{F}$ influences the one-shot key rate (larger $t_{F}$ extends the set of feasible states, thus lowers the key rate) and the acceptance probability of the acceptance test (the larger $t_{F}$, the higher the acceptance probability). We discuss the influence of $t_{F}$ in more detail in Section VI.2, when we compare fixed-length to variable-length key rates.
@@
For the collective attack scenario, this leads to $\epsilon_{\mathrm{AT}}=9.99\times 10^{-11}$, $\epsilon_{\mathrm{PA}}=9.99\times 10^{-11}$ and $\epsilon_{\mathrm{EV}}=1\times 10^{-13}$, resulting in a total security parameter of $\epsilon=1\times 10^{-10}$. In Figure 1a, we examine the key rates for varying $N$ and compare the obtained finite-size key rates with the asymptotic rates for $N\rightarrow\infty$. We fixed the protocol dimension to $d=16$ and the splitting ratio to $r_{S}=1\%$. This means for passive choice settings $r_{S}^{2}$ of the rounds are available for TSUP measurements on both sides and $(1-r_{S})^{2}=0.9801$ of the rounds are measured in ToA by both parties. Given our choice of $m_{W_{1}}=m_{W_{2}}$, $(1-2r_{S})=98\%$ of all rounds are available for key generation. Note that the comparable asymptotic key rates refer to $N\rightarrow\infty$, where the testing fraction can be chosen infinitesimally small. Thus, to ensure a fair comparison, we need to adjust the asymptotic rates by a factor of $0.98$ to take the testing into account. We plot secure key rates for $N\in\{5\times 10^{11},1\times 10^{12},1\times 10^{13},1\times 10^{14},1\times 10^{16},1\times 10^{17}\}$ and observe convergence to the adjusted asymptotic key rate curve, as expected. We obtain non-zero finite-size key rates for block-sizes of $N=5\times 10^{11}$ and above. For $N=10^{17}$, our finite-size results effectively reassemble the asymptotic key rates. We note that the comparably high number of rounds required to obtain a non-zero key is due to the low testing ratio, which we chose to examine asymptotic behaviour (as asymptotic rates only require infinitesimal testing). We will examine higher testing ratios for realistic block-sizes later.
@@
For coherent attacks, the security parameters scale according to Theorem 3. Again, we discuss our choices in detail in Appendix B. Since we chose $r=\frac{1}{1000}=s$, we obtain
@@
Therefore, we aim to keep $\nu_{\mathrm{AT}}$ as large as possible at the cost of lowering the other security parameters more. We choose $\nu_{\mathrm{AT}}=9.98\times 10^{-11}$, $\nu_{\mathrm{PA}}=9.99\times 10^{-14}$, $\nu_{\mathrm{EV}}=1\times 10^{-13}$ such that they add up to a total adjusted security parameter of $\nu=10^{-10}$. The corresponding epsilon parameters can be obtained by scaling according to the security statement in Theorem 3.
@@
In Figure 1b, we plot the obtained secure key rates vs. visibility and examine the asymptotic behavior for large $N$ of the secure key rates against coherent attacks. We fix the dimension again to $d=16$ and the splitting ratio $r_{S}=1\%$ and plot the curves for $N\in\{1\times 10^{16},1\times 10^{17},1\times 10^{18},1\times 10^{19},1\times 10^{20},1\times 10^{22}\}$. We obtain non-zero key rates for $N=10^{17}$ and above and observe convergence to the adjusted asymptotic rate $N\rightarrow\infty$ for large $N$. Again, we note that we chose the low splitting ratio to examine the asymptotic behavior. Since the asymptotic rate does not take testing into account, we again rescale by the corresponding factor of $0.98$. We obtain non-zero key rates for significantly lower $N$ for higher splitting ratios.
@@
Figure 2: Composable secure key rates against i.i.d. collective attacks versus loss for different block-sizes $N$, splitting ratios $r_{s}$ and noise levels, parametrized by the number of solar photons $n_{\mathrm{sol}}$. The system dimension was fixed to $d=16$.
@@
Finally, we examine the behavior with respect to loss for different block-sizes $N$, splitting ratios $r_{s}$, and different noise levels. We use the noise model from Ref. [27] and chose the production rate $c_{\mathrm{prod}}=0.1$ per time-frame for a time-frame length of $T=5.4\times 10^{-9}$. The dark-count rate was assumed to be $100$ per second and the detection efficiency $\eta_{D}=0.90$. In Figure 2, we plot the obtained secure key rates (in bits per pair sent) versus loss. Curves in different shades of blue correspond to $N=10^{10}$ with a splitting ratio to $r_{s}=20\%$, and curves in different shades of red correspond to $N=10^{8}$ with a splitting ratio to $r_{s}=40\%$. Different line styles (solid, dashed, dotted, and dot-dashed) indicate different solar photon rates $n_{\mathrm{sol}}\in\{10^{2},10^{3},10^{4},10^{5}\}$. We observe that for $N=10^{10}$ and $n_{\mathrm{sol}}=10^{2}$, corresponding to almost absolute darkness, we achieve non-zero key rate up to about $51$dB channel attenuation, which is just shy behind the asymptotic rates. For $N=10^{8}$ the tolerable loss decreases to $38.5$dB. While the achievable loss decreases with increasing solar photon rate, in both cases, we still obtain a non-zero key for $n_{\mathrm{sol}}=10^{5}$, corresponding to sunrise [50]. Increasing the splitting ratio $r_{S}$ can improve the noise tolerance even further. This demonstrates that high-dimensional QKD can play a crucial role in overcoming noise in loss regimes present in satellite-based QKD.

### v2

@@ V.2 Formalization of the protocol
[…]
For each round, Alice and Bob may store a measurement outcome $\alpha_{i}$ and $\beta_{i}$ and make a public announcement $a_{i}$ and $b_{i}$, where the measurement outcomes are drawn from the sets $\mathcal{X}_{A}$ and $\mathcal{X}_{B}$ and the announcements from the set $\mathcal{A}_{A}$ and $\mathcal{A}_{B}$ for Alice and Bob, respectively. In our case, these announcements are the choice of measurement, i.e., a number $b$ related to the label of the POVM measurement performed. For the analyzed protocol, $b=1$ corresponds to time-of-arrival measurements, labeled ‘ToA’, while $b=2$ refers to superposition measurements of neighboring time bins, labeled ‘TSUP’. They store their measurement outcomes in their private registers $\bar{A}$ and $\bar{B}$ and their announcements in their public registers $\tilde{A}$ and $\tilde{B}$. Since we perform direct reconciliation, the key map $g$ saves the outcome of the map $g:~\mathcal{X}_{A}\times\mathcal{A}_{A}\times\mathcal{A}_{B}\rightarrow\{0,...,d-1,\perp\}$ to the key register $Z$. In more detail, since the protocol only generates a key when both Alice and Bob measure ToA, the key map for our protocol is trivial and simply copies Alice’s measurement outcome contained in her private register $\bar{A}$ to the key register $Z$. This sets the stage for introducing the formal protocol map. Let $\rho_{AB}$ be Alice’s and Bob’s shared quantum state after the generation and distribution of the high-dimensional entangled state, and let $\rho_{ABE}$ denote Eve’s purification. Furthermore, since Alice’s and Bob’s announcements are public, Eve has access to them. We use the short notation $[E]$ to label Eve’s register, including all available public information. Eve’s goal is to perform some (to us unknown) POVM measurement, aiming to gain as much information as possible about Alice’s measurement outcome. Since Alice and Bob only generate a key if both measure the time of arrival, we obtain the following map, representing their measurement and announcement steps
@@
$\otimes\left[\left(\sqrt{A_{1}^{i}}\otimes\sqrt{B_{1}^{i}}\otimes\sqrt{E^{i}}\right)\rho_{ABE}\left(\sqrt{A_{1}^{i}}\otimes\sqrt{B_{1}^{i}}\otimes\sqrt{E^{i}}\right)^{\dagger}\right],$
@@
$\sigma_{Z[E]}$ $=\mathrm{Tr}_{A\bar{A}B\bar{B}}\left[V\Phi_{\mathrm{M\&A}}V^{\dagger}\right].$ (27)
Nonetheless, since the announcements for this protocol are trivial, they do not give any additional information. Hence, after inserting the expressions for Alice’s and Bob’s key measurement POVM and tracing out the announcement registers, we obtain
@@ VI Results
In this section, we demonstrate our method by deriving composable finite-size key rates for the high-dimensional temporal entanglement protocol and setup analyzed in Ref. [BKPH_2023]. While this is one illustrative example, our method applies to general high-dimensional protocols in a very similar fashion. The underlying data can be found in Ref. [Data].
[…]
First, we cover the fixed-length scenario, which best suits theoretical discussions of achievable ‘one-shot’ key rates, followed by a discussion of variable-length key rates under specific noise models. In all examples given, we use $r_{S}$ to refer to the splitting ratio of the beam splitters in Alice’s and Bob’s labs that passively choose the measurement setting (ToA or TSUP). For example, a splitting ratio of $r_{S}=1\%$ means that $1\%$ of the photons are directed to the TSUP measurement apparatus, while the remaining $99\%$ are measured in the ToA setup. Thus, they obtain $r_{S}^{2}N$ rounds of aligned TSUP measurements and $(1-r_{S})^{2}$ rounds of aligned ToA measurements. Given that 8(d1)8(d-1) different click events are combined into $W_{1}$, we obtain $m_{W_{1}}:=\frac{r_{S}^{2}N}{8(d-1)}$. Additionally, we choose $m_{W_{2}}=m_{W_{1}}$ such that $n:=N(1-r_{S})^{2}-Nr_{s}^{2}=N(1-2r_{s})$, since the click elements required for $W_{2}$ are obtained from ToA measurements.
[…]
Also, note that the leakage in the information reconciliation phase $\mathrm{leak}_{\mathrm{IR}}$ appearing in the key rate formulas given in Section III, is the sum of the error-correction leakage $\mathrm{leak}_{\mathrm{EC}}$ and the length of the hash used for error-verification $\log_{2}\left(\frac{2}{\epsilon_{\mathrm{EV}}}\right)$. If not stated otherwise, we assume ideal error correction.
@@
Figure 2: Composable secure key rates against (a) i.i.d. collective attacks and (b) coherent attacks for different block sizes NN and constant splitting ratio of $1\%$. The system dimension was fixed to $d=16$ and we used the witness $\mathrm{KH1}$.
@@ VI.1 Fixed-length key rates
We start with fixed-length keys, also called ‘one-shot’ key rates, corresponding to key rates usually reported in the literature. Recall that for the definition of the acceptance set (see Theorem 1), one may choose to extend the set of accepted statistics by choosing the parameters $t_{X}>0$ (for each observable $X$). Intuitively, this extends the set of accepted statistics from a point to a (hyper-) rectangle with sides $t_{X}$ for $X\in\Theta$. To simplify notation and to describe the choices for the acceptance test with a single parameter, we measure the chosen $t_{X}$ for the different observables in multiples of their respective $\mu_{X}$ by introducing a factor $t_{F}$, $t_{X}=t_{F}\mu_{X}$. In the whole subsection discussing the results for the fixed-length scenario, we fix $t_{F}=1$. The choice of $t_{F}$ influences the one-shot key rate (larger $t_{F}$ extends the set of feasible states, thus lowering the key rate) and the acceptance probability of the acceptance test (the larger $t_{F}$, the higher the acceptance probability). We discuss the influence of $t_{F}$ in more detail in Section VI.2, when we compare fixed-length to variable-length key rates.
@@
For the collective attack scenario, this leads to $\epsilon_{\mathrm{AT}}=9.99\times 10^{-11}$, $\epsilon_{\mathrm{PA}}=9.99\times 10^{-11}$ and $\epsilon_{\mathrm{EV}}=1\times 10^{-13}$, resulting in a total security parameter of $\epsilon=1\times 10^{-10}$. In Figure 2a, we examine the key rates for varying $N$ and compare the obtained finite-size key rates with the asymptotic rates for $N\rightarrow\infty$. We fixed the protocol dimension to $d=16$ and the splitting ratio to $r_{S}=1\%$. This means for passive choice settings $r_{S}^{2}$ of the rounds are available for TSUP measurements on both sides, and $(1-r_{S})^{2}=0.9801$ of the rounds are measured in ToA by both parties. Given our choice of $m_{W_{1}}=m_{W_{2}}$, $(1-2r_{S})=98\%$ of all rounds are available for key generation. Note that the comparable asymptotic key rates refer to $N\rightarrow\infty$, where the testing fraction can be chosen to be infinitesimally small. Thus, to ensure a fair comparison, we need to adjust the asymptotic rates by a factor of $0.98$ to take the testing into account. We plot secure key rates for $N\in\{5\times 10^{11},1\times 10^{12},1\times 10^{13},1\times 10^{14},1\times 10^{16},1\times 10^{17}\}$ and observe convergence to the adjusted asymptotic key rate curve, as expected. We obtain nonzero finite-size key rates for block sizes of $N=5\times 10^{11}$ and above. For $N=10^{17}$, our finite-size results effectively reassemble the asymptotic key rates. We note that the comparably high number of rounds required to obtain a nonzero key is due to the low testing ratio, which we chose to examine asymptotic behaviour (as asymptotic rates only require infinitesimal testing). We will examine higher testing ratios for realistic block sizes later.
@@
For coherent attacks, the security parameters scale according to Theorem 3. Again, we discuss our parameter choices in detail in Appendix B, and chose $r=\frac{1}{1000}=s$.
@@
We aim to keep $\nu_{\mathrm{AT}}$ as large as possible at the cost of lowering the other security parameters more. We choose $\nu_{\mathrm{AT}}=9.98\times 10^{-11}$, $\nu_{\mathrm{PA}}=9.99\times 10^{-14}$, and $\nu_{\mathrm{EV}}=1\times 10^{-13}$ such that they add up to a total adjusted security parameter of $\nu=10^{-10}$. The corresponding epsilon parameters can be obtained by scaling according to the security statement in Theorem 3.
@@
In Figure 2b, we plot the obtained secure key rates vs. visibility and examine the asymptotic behavior for large $N$ of the secure key rates against coherent attacks. We fix the dimension again to $d=16$ and the splitting ratio $r_{S}=1\%$ and plot the curves for $N\in\{1\times 10^{16},1\times 10^{17},1\times 10^{18},1\times 10^{19},1\times 10^{20},1\times 10^{22}\}$. We obtain nonzero key rates for $N=10^{17}$ and above and observe convergence to the adjusted asymptotic rate $N\rightarrow\infty$ for large $N$. Again, we note that we chose the low splitting ratio to examine the asymptotic behavior. Since the asymptotic rate does not take testing into account, we again rescale by the corresponding factor of $0.98$. We obtain nonzero key rates for significantly lower $N$ for higher splitting ratios.
@@
Figure 3: Composable secure key rates against i.i.d. collective attacks versus loss for different block sizes $N$, splitting ratios $r_{s}$ and noise levels, parametrized by the number of solar photons $n_{\mathrm{sol}}$. The system dimension was fixed to $d=16$.
@@
Next, we examine the behaviour with respect to loss for different block sizes $N$, splitting ratios $r_{s}$, and different noise levels. We use the noise model from Ref. [BKPH_2023] and choose the production rate $c_{\mathrm{prod}}=0.1$ per time frame for a time frame length of $T=5.4\times 10^{-9}$. The dark-count rate was assumed to be $100$ per second, and the detection efficiency $\eta_{D}=0.90$. In Figure 3, we plot the obtained secure key rates (in bits per pair sent) versus loss. Curves in different shades of blue correspond to $N=10^{10}$ with a splitting ratio to $r_{s}=20\%$, and curves in different shades of red correspond to $N=10^{8}$ with a splitting ratio to $r_{s}=40\%$. Different line styles (solid, dashed, dotted, and dot-dashed) indicate different solar photon rates $n_{\mathrm{sol}}\in\{10^{2},10^{3},10^{4},10^{5}\}$. We observe that for $N=10^{10}$ and $n_{\mathrm{sol}}=10^{2}$, corresponding to almost absolute darkness, we achieve a nonzero key rate up to about 5151dB channel attenuation, which is just shy of the asymptotic rates. For $N=10^{8}$, the tolerable loss decreases to 38.538.5dB. While the achievable loss decreases with increasing solar photon rate, in both cases, we still obtain a nonzero key for $n_{\mathrm{sol}}=10^{5}$, corresponding to sunrise [Bulla_2022]. Increasing the splitting ratio $r_{S}$ can improve the noise tolerance even further. This demonstrates that high-dimensional QKD can play a crucial role in overcoming noise in loss regimes present in satellite-based QKD.
As shown in Figure 4, already for $N=10^{10}$, the collective attack key rates (solid blue curve) are only slightly below the asymptotic rates (solid black curve) in terms of tolerable loss and less than half an order of mangitude lower in terms of achievable secure key rate. Additionally, we compare our results with the key rates from Ref. [BKPH_2023] obtained for the same setup, using the same noise model, but relying on a convex optimisation Gauss-Radau method [Araujo_2022]. While Ref. [BKPH_2023] only reports asymptotic rates, similar finite-size corrections are expected, so the comparison of asymptotic curves allows similar conclusions about finite-size results. In Ref. [BKPH_2023], the maximum dimension feasible on a standard computer was $d=8$ (and $d=12$ using a cluster), with the $d=8$ curve (dashed orange) requiring several days of computation. By contrast, our method produces the $d=16$ curve (solid black) within minutes. Already at $d=16$ our method outperforms the numerical $d=8$ result in both key rate and tolerable loss, thereby overcoming the apparent drawback of using $H_{\mathrm{min}}$ instead of the von Neumann entropy. This highlights the clear advantage of our method, which unlocks the use of genuine high dimensions without substantial compromises on the secure key rate.
Figure 4: Comparison of our results with asymptotic behaviour and numerical key rate curves according to Ref. [BKPH_2023]. The splitting ratio was fixed to $r_{s}=20\%$ and the noise level, parametrized by the number of solar photons, was chosen to be $n_{\mathrm{sol}}=10^{2}$. The system dimension varies and is given in the legend.